# GFlowVLM: Enhancing Multi-step Reasoning in Vision-Language Models with Generative Flow Networks

### v1

@@ 4.2 Training Objectives
We adopt three different objective functions of GFlowNets, Variance Trajectory-Balanced (TB) [23], Subtrajectory-Balanced (SubTB) [22], and Detailed-Balanced (DB) [2], to finetune a VLM. We define $z_{t}$ as the state in the trajectory sequence that includes both $a_{t}$, $s_{t}$, and $o_{t}$.
@@ 4.2.1 Variance Trajectory Balanced (Var-TB) Loss
The Trajectory-Balanced (TB) objective ensures that the probability of generating a complete trajectory $\tau=(z_{0}\rightarrow z_{1}\rightarrow\cdots\rightarrow z_{n}\rightarrow x)$ is proportional to the reward $R(x)$. This objective is given by:
$\mathcal{L}_{\mathrm{VarTB}}(\boldsymbol{\tau};\theta)=\frac{1}{N}\sum_{i=1}^{N}\bigg{(}\mathrm{\zeta}(\tau_{i};\theta)-\mathbb{E}_{\boldsymbol{\tau}}\big{[}\mathrm{\zeta}(\boldsymbol{\tau};\theta)\big{]}\bigg{)}^{2},$ (6)
where $\zeta(\cdot)$ is the estimated initial flow (see Eq. 13 for details) and $N$ represents the number of sampled trajectories. The details are available in the Supplementary material. This loss ensures that the high-reward trajectories are sampled more frequently by the policy.
@@ 4.2.2 Subtrajectory Balanced (SubTB) Loss
$\begin{split}\mathcal{L}_{\mathrm{SubTB}}(z_{0:m},g;\theta)=\sum_{0\leq i<j\leq m}\\ \Bigg{(}\log\frac{R(z_{0:i}\top)\prod_{k=i+1}^{j}P_{F}(z_{k}|z_{0:k-1},g;\theta)P_{F}(\top|z_{0:j},g;\theta)}{R(z_{0:i}\top)P_{F}(\top|z_{0:i},g;\theta)}\bigg{)}^{2}\end{split}$ (7)
where $\top$ is the [DONE] symbol, and process continues until [DONE] symbol $\top$ is generated similar to [8]. This loss penalizes discrepancies in local transitions and ensures that all subsegments of a trajectory follow the correct balance conditions, reducing variance in smaller parts of the trajectory.
@@ 4.2.3 Detailed Balanced (DB) Loss
Input: An environment env, an initial VLM with parameters $\theta_{0}$, a CoT reasoning scaling factor $\lambda$, maximum episode length $T$, number of tasks $W$, number of collected trajectories per task $K$.
@@ 5 Experiments
We evaluate the performance of GFlowVLM on three distinct tasks that require multi-step visual-language reasoning. The Numberline and Blackjack tasks assess GFlowVLM’s arithmetic reasoning capabilities with maximum steps set as 10, while Alfworld focuses on decision-making tasks that demand visual-semantic understanding with max steps set as 35 in a sequence. We mainly compare the performance with RL4VLM [41] and SFT methods. For fair comparison, we use the same base VLM of LLAVA-v1.6-Mistral-7B [19]. We conduct 4 independent runs with different random seeds, reporting mean and standard deviation. Episode success rate measures reasoning performance across tasks, while the diversity metric (Div@N[40] quantifies unique correct solutions across N samples. The minimum for Div@N is 1. GFlowNets’ reward function is modified for Numberline and Blackjack as outlined in Sec. 5.2.
@@ 5.1 Baselines
We employ two versions of SFT in our baseline: SFT-w/o-[DONE] and SFT-w/-[DONE]. SFT-w/o-DONE uses the same GPT-4o dataset as in [41]. For SFT-w/-DONE, we include the [DONE] action in the training inputs and add correct examples where outputs explicitly contain [DONE]. We fine-tune the LLaVA-1.6-7B model on this dataset for 1 epoch using the official script. To ensure consistency, downstream GFlowNets training for both SubTB and DB losses starts from the same SFT checkpoint that includes [DONE].
@@
5.2 Environments
Numberline. This task involves moving a current number “Current: $y_{t}$” to a “Target: $t$”. The agent’s goal is to align $y_{t}$ with $t$ by outputting an action $a_{t}$ from the discrete set {“+”, “-”, [DONE](if applicable)}. In-distribution examples include numbers from 0 to 5, and OOD examples range from 10 to 50. We revise the reward function to replace the original discrete rewards of -1, 0, and 1 with non-negative values as follows: $R(\text{target},\text{current})=\frac{c}{|\text{target}-\text{current}|+1}$, where $c$ is a scaling constant set to 100. This reward incentivizes the model to bring the current number closer to the target, progressively increasing the reward as the gap decreases. For fair comparison, we run RL4VLM [41] with revised reward structure.
Blackjack. The Blackjack task requires VLM to reason with visual information and adapt to stochastic outcomes. The agent aims to win by selecting an action $a_{t}$ from {“stand”, “hit”, [DONE](if applicable)}. We revise the reward function to replace the original discrete rewards of -1, 0, and 1 with non-negative values as follows: $R(s_{t},a_{t})=\max(1\times 10^{-10},(r(s_{t},a_{t})+1)\times 10)$, where $r(s_{t},a_{t})$ represents the environment’s original reward for state $s_{t}$ and action $a_{t}$. This scales the rewards and ensures they are strictly non-negative. For fair comparison, we run RL4VLM [41] with revised reward structure.
ALFWorld ALFWorld [31] is an embodied AI environment combining a text-based interactive setup with a vision-language planning dataset. ALFWorld has a state-dependent action space $\mathcal{A}_{t}$; Our prompt instructs the VLM to choose from the admissible actions $\mathcal{A}_{t}$, and we evaluate out-of-distribution (OOD) performance using a test set of previously unseen scenes. We use the same non-negative reward function as used in [41], making it suitable for Var-TB and SubTB losses. However, since it lacks dense rewards for every transition, DB does not perform effectively.
@@ 5.2 Environments
6 Results Analysis
Diverse Solutions
Our method generates more diverse solutions than other baselines. In ALFWorld, GFlowNets achieve 25% and 33% higher diversity than RL4VLM and SFT ( Tab. 4), as measured by diversity metric, capturing a wider range of plausible solutions, offering a distinct advantage in scenarios that benefit from broader strategy exploration.
Improved Generalization on OOD samples
@@ 6 Results Analysis
Figure 3: The training curves of in-distribution episode success rates (%) of different methods across three tasks. For the Numberline and BlackJack, RL4VLM has been trained with original reward function, while the GFlowVLM with various loss functions have been trained with revised reward function. This is because RL4VLM serves a strong baseline when trained with original rewards. Since, we did not revise the rewards for ALFWorld, we used the same reward function across all the methods. These models are trained with On-policy sampling method.
@@
Since GFlowNets allow for off-policy [40, 24, 13] along with on-policy learning unlike PPO [29], we adopt an off-policy data generation approach to evaluate the impact of using more accurate trajectories during training.
  • Numberline: During training, if the on-policy trajectory produced by the model fails to correctly move the current number towards the target, we add to buffer an additional trajectory with an off-policy ground-truth trajectory. This ensures that the model learns from successful trajectories, even when its own predictions deviate from the correct trajectory. By integrating these ground-truth trajectories, we aim to improve the model’s ability to generalize to new instances by leveraging more accurate training data.
  • Blackjack: Due to the stochastic nature of this task, deterministic ground-truth trajectories are unavailable. Instead, we generate high-quality off-policy trajectories using a rule-based approach: the agent “stands” if the hand value is 17 or higher; otherwise, it “hits”. This aligns with basic Blackjack strategy, where drawing on a hand of 17 or more risks exceeding 21, balancing improvement chances with busting risk.
@@
Tab. 3 shows results for NumberLine (NL), out-of-distribution NumberLine (NL-OOD), and BlackJack (BJ) tasks. GFlowVLM with Var-TB, SubTB, and DB, demonstrate improvements with offline data , averaging a 36.2% performance increase over online-only approaches. These results indicate that each loss function benefits from off-policy high-quality data, leveraging both correct and incorrect solutions to enhance performance, even in challenging OOD scenarios.
Training Efficiency
@@
As shown in  Fig. 3, GFlowNets converge faster than RL4VLM on NumberLine, Blackjack, and ALFWorld, reaching optimal performance with significantly fewer environment steps—about 10,000 fewer than RL4VLM. This efficiency reduces training time and computational demands, supporting scalability for complex reasoning tasks.
Comparisons with different loss functions
@@
As shown in the training curves, all three loss functions—DB, Var-TB, and SubTB-converge at a similar rate. DB, which requires dense rewards for every transition since it utilize transitions as input, demonstrates the best generalization, as evidenced in  Tab. 3, in the NumberLine and Blackjack tasks. Both SubTB and TB achieve comparable performance in terms of in-distribution and OOD generalization, making them equally effective for a wide range of reasoning tasks.
SFT Initialization for SubTB and DB Losses
@@
The termination probability $P_{F}(\top|\cdot)$ in DB and SubTB losses estimates the modified flow in GFlowNet [27], which Var-TB lacks. To enable VLMs to accurately model this for DB and SubTB, we first apply SFT on correctly completed trajectories before fine tuning with GFlowNets. As shown in Tab. 3, SFT initialization significantly boosts SubTB and DB performance on the NumberLine and Blackjack tasks. Without SFT, both losses perform poorly, especially on NL and BJ tasks. With SFT, SubTB and DB improve by 50% and 36% for NumberLine and by 107% and 103% for Blackjack, largely due to better estimation of terminal probability $P_{F}(\top|\cdot)$.
Markovian and non-Markovian assumptions
@@
GFlowVLM outperforms RL4VLM in Non-Markovian settings, excelling in complex, long-horizon tasks. In ALFWorld (Tab. 4), GFlowVLM achieves higher average performance by 18%, OOD robustness by 100%, and diversity by 27%. It also achieves better success rates in gym tasks (Tab. 3), where history aids decision-making, underscoring GFlowNets’ advantage over PPO-based methods.
@@ 7 Conclusion, Limitation, Future Works
We introduced a novel framework using GFlowNets to enhance structured reasoning in Vision Language Models (VLMs), to capture relationships among reasoning steps for improved generalization. Unlike traditional methods like Supervised Fine-Tuning and Proximal Policy Optimization, which are limited by certain assumptions, our approach supports complex, long-term reasoning tasks. Experiments in card games and embodied planning showed enhanced training efficiency, diversity, and generalization. We focus on a single-agent task setting, leaving multi-agent task and alternative prompting methods as future directions. Limited computational resources led us to use small sized VLMs, but larger models may further benefit from GFlowNets.
Appendix A Environments
A.1 ALFWorld
@@
ALFWorld [31] is an embodied AI environment combining a text-based interactive setup with a vision-language planning dataset. It includes six goal-conditioned tasks: “Pick & Place”, “Examine in Light”, “Clean & Place”, “Heat & Place”, “Cool & Place”, and “Pick Two & Place”. The agent must plan and act based on visual cues and textual instructions (e.g., “go to shelf 1”) that specify the task. Unlike gym_cards, where all states share the same action space, ALFWorld has a state-dependent action space $\mathcal{A}_{t}$; actions are context-dependent (e.g., “put some pillows on the armchair”, the agent can only place a pillow after picking it up). Our prompt instructs the VLM to choose from the admissible actions $\mathcal{A}_{t}$, and we evaluate Out-Of-Distribution (OOD) performance using a test set of previously unseen scenes (see detailed prompt templates in  Tab. 9 and  Tab. 10). We use the same non-negative components of the reward function used in [41], which includes sub-goal and goal rewards: $r(s_{t},a_{t},s_{t+1}|g_{task})=50*1\{s_{t+1}=g_{task}\}+1\{s_{t+1}=g_{task}\}$. We do not include the negative component of the reward function represented as $-1\{a_{t}\notin\mathcal{A}_{t}(s_{t})\}$ in [41], since the actions are always selected from the admissible actions provided in the input prompt $p_{t}$. The rewards are non-negative, making it suitable for Var-TB and SubTB losses. However, since it lacks dense rewards for every transition, we didn’t GFlowVLM with DB loss.
A.2 NumberLine

### v2

@@ 4.2 Training Objectives
We adopt three different objective functions of GFlowNets, Variance Trajectory-Balanced (TB) [23], Subtrajectory-Balanced (SubTB) [22], and Detailed-Balanced (DB) [2], to finetune a VLM. We define $z_{t}$ as the state in the trajectory sequence that includes both $a_{t}$, $s_{t}$, and $o_{t}$. See the Sec. C of the Supplementary Material for more details.
@@ 4.2.1 Variance Trajectory Balanced (Var-TB) Loss
The Trajectory-Balanced (TB) objective ensures that the probability of generating a complete trajectory $\tau=(z_{0}\rightarrow z_{1}\rightarrow\cdots\rightarrow z_{n}=x)$ is proportional to the reward $R(x)$. This objective is given by:
$\mathcal{L}_{\mathrm{VarTB}}(\boldsymbol{\tau};\theta)=\frac{1}{K}\sum_{k=1}^{K}\bigg{(}\mathrm{\zeta}(\tau_{k};\theta)-\mathbb{E}_{\boldsymbol{\tau}}\big{[}\mathrm{\zeta}(\boldsymbol{\tau};\theta)\big{]}\bigg{)}^{2},$ (6)
where $\zeta(\cdot)$ is the estimated initial flow (see  Eq. 13 in Supplementary Material for details), $\tau_{k}$ is $k^{th}$ sampled trajectory during training, and $K$ represents the total number of trajectories. This loss ensures that the high-reward trajectories are sampled more frequently by the policy.
@@ 4.2.2 Subtrajectory Balanced (SubTB) Loss
$\begin{split}\mathcal{L}_{\mathrm{SubTB}}(z_{0:m},g;\theta)=\sum_{0\leq i<j\leq m}\\ \Bigg{(}\log\frac{R(z_{0:i}\top)\prod_{k=i+1}^{j}P_{F}(z_{k}|z_{0:k-1},g;\theta)P_{F}(\top|z_{0:j},g;\theta)}{R(z_{0:j}\top)P_{F}(\top|z_{0:i},g;\theta)}\bigg{)}^{2}\end{split}$ (7)
where $i$ and $j$ are two time steps along a subtrajectory, and $\top$ is the [DONE] symbol to terminate the trajectory. $\top$ is generated similar to [8]. The reward $R(z_{0:i}\top)$ is computed as a cumulative reward given by the environment from step 0 to step $i$. This loss penalizes discrepancies in local transitions and ensures that all subsegments of a trajectory follow the correct balance conditions, reducing variance in smaller parts of the trajectory.
@@ 4.2.3 Detailed Balanced (DB) Loss
Input: An environment env, an initial VLM with parameters $\theta_{0}$, a CoT reasoning scaling factor as $\lambda$, maximum episode length as $T$, number of tasks as $W$, number of collected trajectories per task as $K$.
@@ 5 Experiments
We evaluate the performance of GFlowVLM on three distinct tasks that require multi-step visual-language reasoning. The Numberline and Blackjack tasks assess GFlowVLM’s arithmetic reasoning capabilities with maximum steps set as 10, while Alfworld focuses on decision-making tasks that demand visual-semantic understanding with max steps set as 35 in a sequence. We mainly compare the performance with RL4VLM [41] and SFT methods. For fair comparison, we use the same base VLM of LLAVA-v1.6-Mistral-7B [19]. We conduct 4 independent runs with different random seeds, reporting mean and standard deviation. Episode success rate measures reasoning performance across tasks, while the diversity metric (Div@N[40] quantifies unique correct solutions across N samples. The minimum for Div@N is 1. NumberLine and Blackjack have discrete negative rewards, but GFlowNets inherently do not support negative rewards (Sec. 3). Thus, we apply reward shaping as outlined in Sec. 5.2 on these two tasks.
@@ 5.1 Baselines
We employ two versions of SFT in our baseline: SFT-w/o-[DONE] and SFT-w/-[DONE]. SFT-w/o-[DONE] uses the same GPT-4o dataset as in [41]. For SFT-w/-[DONE], we include the [DONE] action in the training inputs and add correct examples where outputs explicitly contain [DONE]. We fine-tune the LLaVA-1.6-7B model on this dataset for 1 epoch using the official script. To ensure consistency, downstream GFlowNets training for both SubTB and DB losses starts from the same SFT checkpoint that includes [DONE].
@@
5.2 Environments
Numberline. This task involves moving a current number “current: $y_{t}$” to a “target: $c$”. The agent’s goal is to align $y_{t}$ with $c$ by outputting an action $a_{t}$ from the discrete set {“+”, “-”, [DONE](if applicable)}. In-distribution examples include numbers from 0 to 5, and OOD examples range from 10 to 50. We revise the reward function to replace the original discrete rewards of -1, 0, and 1 with non-negative values as follows: $R(x)=R(c,y_{t})=\frac{l}{|c-y_{t}|+1}$, where $l$ is a scaling constant set to 100. This reward incentivizes the model to bring the current number closer to the target, progressively increasing the reward as the gap decreases. For fair comparison, we run RL4VLM [41] with revised reward structure.
Blackjack. The Blackjack task requires VLM to reason with visual information and adapt to stochastic outcomes. The agent aims to win by selecting an action $a_{t}$ from {“stand”, “hit”, [DONE](if applicable)}. We revise the reward function to replace the original discrete rewards of -1, 0, and 1 with non-negative values as follows: $R(x)=\max(1\times 10^{-10},(r(x)+1)\times 10)$, where $r(x)$ represents the environment’s original reward for terminal state $x$. This scales the rewards and ensures they are strictly non-negative. For fair comparison, we run RL4VLM [41] with revised reward structure.
ALFWorld. ALFWorld [31] is an embodied AI environment combining a text-based interactive setup with a vision-language planning dataset. ALFWorld has a state-dependent action space $\mathcal{A}_{t}$; Our prompt instructs the VLM to choose from the admissible actions $\mathcal{A}_{t}$, and we evaluate out-of-distribution (OOD) performance using a test set of previously unseen scenes. We use the same non-negative reward function as used in [41], making it suitable for Var-TB and SubTB losses. However, since it lacks dense rewards for every transition, DB does not perform effectively.
@@ 5.2 Environments
6 Results Analysis
Diverse Solutions
Our method generates more diverse solutions than other baselines. In ALFWorld, GFlowNets achieve 25% and 33% higher diversity than RL4VLM and SFT (Tab. 4), as measured by diversity metric, capturing a wider range of plausible solutions, offering a distinct advantage in scenarios that benefit from broader strategy exploration. In contrast to PPO, which optimizes a single best policy for long-term planning, GFlowNet finds multiple diverse high-reward solutions, making it better suited for structured generation (see Section E for detailed qualitative results).
Improved Generalization on OOD samples
[…]
@@ 6 Results Analysis
Since GFlowNets allow for off-policy [40, 24, 13] along with on-policy learning unlike PPO [29], we adopt an off-policy data generation approach to evaluate the impact of using more accurate trajectories during training (see Section D.1 for the details of data generation). Tab. 3 shows results for NumberLine (NL), out-of-distribution NumberLine (NL-OOD), and BlackJack (BJ) tasks. GFlowVLM with Var-TB, SubTB, and DB, demonstrate improvements with offline data, averaging a 36.2% performance increase over online-only approaches. These results indicate that each loss function benefits from off-policy high-quality data, leveraging both correct and incorrect solutions to enhance performance, even in challenging OOD scenarios.
@@
Training Efficiency. As shown in  Fig. 3, GFlowNets converge faster than RL4VLM on NumberLine, Blackjack, and ALFWorld, reaching optimal performance with significantly fewer environment steps—about 10,000 fewer than RL4VLM. This efficiency reduces training time and computational demands, supporting scalability for complex reasoning tasks.
@@
Comparisons with different loss functions. As shown in the training curves, all three loss functions—DB, Var-TB, and SubTB-converge at a similar rate. DB, which requires dense rewards for every transition since it utilize transitions as input, demonstrates the best generalization, as evidenced in  Tab. 3, in the NumberLine and Blackjack tasks. Both SubTB and TB achieve comparable performance in terms of in-distribution and OOD generalization, making them equally effective for a wide range of reasoning tasks.
@@
SFT Initialization for SubTB and DB Losses. The termination probability $P_{F}(\top|\cdot)$ in DB and SubTB losses estimates the modified flow in GFlowNet [27], which Var-TB lacks. To enable VLMs to accurately model this for DB and SubTB, we first apply SFT on correctly completed trajectories before fine tuning with GFlowNets. As shown in Tab. 3, SFT initialization significantly boosts SubTB and DB performance on the NumberLine and Blackjack tasks. Without SFT, both losses perform poorly, especially on NL and BJ tasks. With SFT, SubTB and DB improve by 50% and 36% for NumberLine and by 107% and 103% for Blackjack, largely due to better estimation of terminal probability $P_{F}(\top|\cdot)$.
@@
Markovian and non-Markovian assumptions. GFlowVLM outperforms RL4VLM in non-Markovian settings, excelling in complex, long-horizon tasks. In ALFWorld (Tab. 4), GFlowVLM achieves higher average performance by 18%, OOD robustness by 100%, and diversity by 27%. It also achieves better success rates in gym tasks (Tab. 3), where history aids decision-making, underscoring GFlowNets’ advantage over PPO-based methods.
@@ 7 Conclusion, Limitation, Future Works
We introduce a novel framework using GFlowNets to enhance structured reasoning in VLMs, to capture relationships among reasoning steps for improved generalization. Unlike traditional methods like SFT and PPO, which are limited by certain assumptions, our approach supports complex, long-term reasoning tasks. Experiments in card games and embodied planning showed enhanced training efficiency, diversity, and generalization. We focus on a single-agent task setting, leaving multi-agent task and alternative prompting methods as future directions. Limited computational resources led us to use small-sized VLMs, but larger models may further benefit from GFlowNets.
Appendix A Preliminaries
A.1 GFlowNets
We summarize the necessary preliminaries of GflowNets and encourage readers to refer to [2] for deeper understanding. In a directed acyclic graph $\mathcal{G}=(\mathcal{S},\mathcal{A})$ with states $\mathcal{S}$ and directed actions $\mathcal{A}$, a complete trajectory is any trajectory starting in initial state $s_{0}$ and ending in terminal state $x\in X$ where $X\subset\mathcal{S}$. There is a unique initial state $s_{0}\in S$ with no parents. States with no children are called terminal, and the set of terminal states is denoted by $\mathcal{X}$. A trajectory $\tau=(s_{0}\to\ldots\to s_{n}=x)$ represents a complete sequence ending in a terminal state $x\in\mathcal{X}$ where each $(s_{t}\to s_{t+1})$ is an action. The trajectory flow $F:\mathcal{T}\rightarrow\mathbb{R}_{\geq 0}$ defines flows over trajectories, with state flow $F(s)=\sum_{s\in\tau}F(\tau)$ and with edge flow $F(s\to s^{\prime})=\sum_{\tau=(\ldots\to s\to s^{\prime}\to\ldots)}F(\tau)$. The trajectory flow $F$ is Markovian if there exist action distributions $P_{F}(\cdot|s)$ over the children of each non-terminal state $s$.
A.1.1 Forward and Backward Policies
A forward policy $P_{F}(\cdot|s)$, often parametrized by a neural network, induces a distribution over trajectories and a marginal distribution over the children of every non-terminal state $s\in S$, with probabilities given by: $P_{F}(\tau)=P_{F}(s_{0}\to\ldots\to s_{n})=\prod_{t=0}^{n-1}P_{F}(s_{t+1}|s_{t})\quad\forall\tau\in\mathcal{T}$. The distribution over complete trajectories that arises from a forward policy satisfies a Markov property. The forward policy can then be used to sample terminal states $x\in X$ by starting at state $s_{0}$ and iteratively sampling actions from $P_{F}$. A backward policy $P_{B}(\tau)=P_{B}(s_{n}\to\ldots\to s_{0})=\prod_{t=0}^{n-1}P_{B}(s_{t}|s_{t+1})\quad\forall\tau\in\mathcal{T}$. If $F$ is markovian flow, then $P_{F}$ and $P_{B}$ can be computed in terms of state and edge flows as: $P_{F}(s^{\prime}|s)=\frac{F(s\to s^{\prime})}{F(s)}$ and $P_{B}(s|s^{\prime})=\frac{F(s\to s^{\prime})}{F(s^{\prime})}$.
Given a non-negative reward function $R:\mathcal{X}\rightarrow\mathbb{R}_{\geq 0}$, GFlowNets aim to learn a policy such that the probability of sampling a state $x\in\mathcal{X}$ is proportional to $R(x)$. The marginal likelihood of sampling a state $x\in X$ is the sum of likelihoods of all complete trajectories that terminate at $x$. If the objective function is globally minimized, then the likelihood of terminating at state $x$ is proportional to $R(x)$. Formally, the learning problem solved by GFlowNets is to estimate a policy $P_{F}$ over trajectories such that there exists a normalizing constant $Z$ satisfying: $R(x)=Z\sum_{\tau=(s_{0}\rightarrow\ldots\rightarrow s_{n}=x)}P_{F}(\tau)\quad\forall x\in\mathcal{X}$, where $Z=F(s_{0})=\sum_{\tau\in\mathcal{T}}F(\tau)$ is total flow at the initial state, and $\tau\in\mathcal{T}$ is the trajectory.
A.2 Motivating Example
Figure 4: Overview of the prediction of diverse sequence using GFlowVLMs as compared to PPO for AlfWorld scenarios. The model takes the image of sequence and prompt as input, and generates the next number of sequence by implicitly modeling the causality.
We include a practical example from ALFWorld demonstrating how GFlowVLM can be applied to embodied AI tasks in Fig. 4. The agent is presented with a visual observation of a simulated household environment and a high-level goal in natural language, such as ”Put keychain in ottoman.” The task requires the agent to generate a valid sequence of actions (e.g., open drawer $\rightarrow$ take keychain from drawer $\rightarrow$ close the opened drawer $\rightarrow$ go to ottomann $\rightarrow$ place keychain in ottoman). Importantly, there are multiple valid plans that achieve the same goal, with subtle causal constraints (e.g., the drawer must be open before taking the keychain from it, and objects must be picked up before being moved). We observe that models trained with PPO tend to converge on the most common or shortest path, while GFlowVLM generates a more diverse set of valid action sequences, reflecting a richer understanding of the causal structure of the environment. This example demonstrates GFlowNets’ strength in reasoning over multimodal inputs and learning structured, stochastic policies that preserve functional diversity.
Appendix B Environments
B.1 ALFWorld
@@
ALFWorld [31] is an embodied AI environment combining a text-based interactive setup with a vision-language planning dataset. It includes six goal-conditioned tasks: “Pick & Place”, “Examine in Light”, “Clean & Place”, “Heat & Place”, “Cool & Place”, and “Pick Two & Place”. The agent must plan and act based on visual cues and textual instructions (e.g., “go to shelf 1”) that specify the task. Unlike gym_cards, where all states share the same action space, ALFWorld has a state-dependent action space $\mathcal{A}_{t}$; actions are context-dependent (e.g., “put some pillows on the armchair”, the agent can only place a pillow after picking it up). Our prompt instructs the VLM to choose from the admissible actions $\mathcal{A}_{t}$, and we evaluate Out-Of-Distribution (OOD) performance using a test set of previously unseen scenes (see detailed prompt templates in  Tab. 9 and  Tab. 10). We use the same non-negative components of the reward function used in [41], which includes sub-goal and goal rewards: $r(s_{t},a_{t},s_{t+1}|g_{task})=50*1\{s_{t+1}=g_{task}\}+1\{s_{t+1}=g_{task}\}$. We do not include the negative component of the reward function represented as $-1\{a_{t}\notin\mathcal{A}_{t}(s_{t})\}$ in [41], since the actions are always selected from the admissible actions provided in the input prompt $p_{t}$. The rewards are non-negative, making it suitable for Var-TB and SubTB losses. However, since it lacks dense rewards for every transition, we didn’t use GFlowVLM with DB loss on this task.
B.2 NumberLine